OpenStudy (anonymous):

Find the quotient 5 --- 43 ÷ 25 --- 86

sam (.sam.):

\[\frac{5}{43}\div\frac{25}{86}\] \[\frac{5}{43}\times\frac{86}{25}\] \[\frac{2}{5}\]